### v1

@@ Abstract
The synergy between deep learning models and traditional automation tools plays a pivotal role in developing robust neural theorem provers (NTPs). However, for proof synthesis with LLMs, previous work applies automation tools either only when the model explicitly calls the method, or only at a single granularity level, failing to fully exploit the power of built-in tactics and off-the-shelf automated theorem provers. In this work, we propose ProofAug, a novel theorem proving method that enjoys superior sample efficiency through equipping proof-generation LLMs with automation methods in different granularities via fine-grained structure analysis of model-generated proof proposals. Furthermore, ProofAug serves as a versatile plug-and-play module that seamlessly integrates with any tree-search algorithm, enabling our construction of an efficient recursive proving (ERP) module to further enhance performance. The superiority of our method is validated on the miniF2F-test benchmark using the open-source deepseek-math-7b-base model and the Isabelle proof assistant. Notably, by additionally employing a mixed prompting strategy, we achieve a cumulative pass rate of 66.0% after curation of the dataset (61.9% for the original version), setting a new SOTA across all proof languages with a total sample budget of only 2100. Our code is available at https://github.com/haoxiongliu/ProofAug.
@@ 1 Introduction
Automated theorem proving is a field that not only appeals to mathematicians seeking efficient formalization and proof automation for mathematical theorems, but also finds significant applications in real-world industries such as integrated circuit design and software engineering (Bibel, 2013; Schumann, 2013). Early applications of machine learning methods on automated theorem proving focus on tasks such as premise selection (Irving et al., 2016) and proof search (Loos et al., 2017) in SMT solvers or resolution/superposition provers, often interleaved with handcrafted heuristics. Recently, the rapid progress in mathematical reasoning with large language models (LLMs) has awoken a flourishing interest for directly synthesizing formal proofs using generative language models.
@@
There are primarily two paradigms for direct proof synthesis: the proof-step generation paradigm and the whole-proof generation paradigm, each corresponding to a distinct proof style. For proof assistants where proofs are mainly in the procedural style (or tactic style), such as MetaMath (Megill & Wheeler, 2019) and HOL Light (Harrison, 2009), previous work adopts the proof-step generation paradigm (Polu & Sutskever, 2020; Lample et al., 2022; Han et al., 2022; Polu et al., 2022), where in each iteration the model is requested to generate one tactic to be applied in the interactive theorem prover (ITP). The ITP provides information about the updated proof state, which will be included in the prompt for the generation of next tactic. Proof-step generation methods typically use tree-search algorithms to guide the proving process, inspired by the success of AlphaGo (Silver et al., 2016). In contrast, for proof systems that support declarative (or term-style) proofs such as Isabelle and Lean, computation-efficient whole-proof generation methods become possible. Nevertheless, the sample efficiency under this paradigm is not free. Any incorrect application of a proof method bridging the gap between two proof terms (or goals/conjectures) will cause the proof to fail to pass the verification check of the ITP.
@@
To avoid this issue, the Draft, Sketch, and Prove (DSP) framework (Jiang et al., 2023) proposes to let the model first write a formal proof sketch (Wiedijk, 2004) according to an informal proof draft, then offload the proofs of intermediate conjectures to off-the-shelf ATPs [^1] For simplicity, we use the abbreviation ATP to refer to both SMT solvers and automated theorem provers rooted in the tradition of TPTP (Sutcliffe, 2009; Blanchette et al., 2013). This is also the practice in Jiang et al. (2023). and heuristic proof methods. Many subsequent works (Wang et al., 2023a; Zhao et al., 2023; Zheng et al., 2024; Zhao et al., 2024) also follow the practice of DSP when designing their few-shot prompts: The proof methods in the original proofs of the demonstration examples are replaced with a placeholder, which signifies that ATPs are responsible for completing the task. We observe that while this approach can significantly enhance performance compared to direct full-proof generation, it may also result in proof failures in unintended ways:
@@
  • (Complicated Draft) Due to the mismatch between informal proof and formal system, there are conjectures that despite being easy to be solved by ATPs, humans need to write complicated informal proofs for them. The unnecessary intermediate steps could be translated into formal conjectures that are inappropriate in syntax or even harder than the original theorem.
@@
Figure 1: Illustration of our ProofAug method by an example. Each box in the flowchart corresponds to a proof step. (a) The initial proof encounters an error when proving Conjecture 1-0. (b) We replace the original proof of Conjecture 1-0 with a sorry and continue, until a syntactic error occurs at Conjecture 1-1. (c) The proof of Conjecture 1 is replaced by sorry and we obtain a semi-proof with two sorrys. (d) Automation tools are called to prove Conjecture 0 and Conjecture 1, but failing at the latter one. (e) The ERP module (introduced in Section 3.3) sends a completion request at Conjecture 1. (f) The generated ERP proof still fails at Conjecture 1. (g) We finally resort to a more coarse level of proof and successfully find a proof with automation tools. Refer to Algorithm 2 for a pseudo-code description of ProofAug.
@@
In this paper, we propose to address this dilemma through performing fine-grained structural analysis on the proof proposals generated by LLMs, based on the observation that regardless of correctness, full proofs contain rich structural information from which valid semi-proofs [^3] By semi-proofs, we refer to proofs that possibly contain symbols indicating pending proof, such as sorry in Isabelle. of varying granularities can be derived. Specifically, given a theorem statement, instead of instructing the model to generate a proof sketch as previous work does, we choose to start from generating a full proof to suppress the ‘Hard Conjectures’ issue in the first place. For the generated proof proposal, we first find its corresponding maximal compatible semi-proof (MCSP) that maintains most structural information yet can pass the check of ITP. Then, departing from it, we recursively resort to a more coarse semi-proof whenever ATPs fail to fill in some gap in the current one. This procedure solves the ‘Complicated Draft’ issue, thus significantly improving the sample efficiency. Figure 1 shows an example flow of our method.
@@
We also demonstrate that the application of proof structure analysis is not restricted to single-pass generation. In fact, ProofAug can be incorporated into any tree-search algorithm to expand the candidate node set at each iteration. This will become evident in our unified view of theorem proving introduced in Section 2.1. In Section 3.3, we illustrate its application in enabling a sample-efficient approach to recursive theorem proving (Wang et al., 2024).
@@
We list our contributions with corresponding key evaluation results using below. All experiments use the open-source deepseek-math-7b-base model (Shao et al., 2024).
@@
  • We propose a novel neural theorem proving method, ProofAug, based on fine-grained proof structure analysis. On miniF2F-test (Zheng et al., 2021), ProofAug achieves a pass rate of 52.5% with no more than 100 queries to the model. To the best of our knowledge, no previous work reports a higher pass rate with equal or fewer times of queries. Ablation results show that when compared with the DSP baseline, the improvement brought by ProofAug is 7.8%, 4.1% and 3.3% under sample budgets of 1, 10 and 100, respectively.
@@
  • ProofAug can be incorporated into any tree-search algorithms under our unified view of existing generative LM based theorem proving methods. We showcase that when integrated with our efficient recursive proving (ERP) module, 0-shot ProofAug performance improves from 54.5% to 56.1% with a limit of 500 queries per problem. As a comparison, RMaxTS (Xin et al., 2024b) only shows $<0.2\%$ advantage over naive single-pass retrying under a sample budget of 3200.
@@
  • Moreover, using a mixture of strategies with different promptings, we achieve a total pass rate of 61.9% within 1700 model queries. After curating the miniF2F-test data, we manage to prove 66.0% of the theorems with no more than 2400 requests, setting a new state of the art across all proof languages on this benchmark.
@@ 2.1 A unified view of theorem proving with generative language models
  • $\mathcal{A}\subset\Sigma^{*}$ is the set of proof steps. There is also a parser $\operatorname{Parse}:\Sigma^{*}\rightarrow A^{*}$ that transforms a string into a sequence of proof steps.
@@
For any string $x\in\Sigma^{*}$, we use $s_{x}$ to denote $T(s_{0},x)$, where we have lifted the notation $T$ to express consecutive state transitions, i.e., for $s\in\mathcal{S}$, $T(s,x)=T(\cdot,\mathbf{a}[n])\circ T(\cdot,\mathbf{a}[n-1])\circ\cdots\circ T(\cdot,\mathbf{a}[1])(s)$ if $\mathrm{Parse}(x)=(\mathbf{a}[1],\mathbf{a}[2],\cdots,\mathbf{a}[n])$.
@@
2.2 Isabelle/Isar Proof Language
Although the basic idea of our method in general fits in any ITP [^4] See Appendix C for a detailed explanation of why our method is generic across different systems., the implementation details vary depending on specific features of the proof assistant. In this section we introduce some basic knowledge and unique features that ease the implementation of our method in Isabelle/Isar (Wenzel et al., 2004), the proof language used in this work.
@@ 3 Method
Algorithm 2 describes our method, ProofAug. It mainly consists of three parts: finding the Maximal Compatible Semi-Proof (MCSP), proof augmentation, and an optional Efficient Recursive Proving (ERP) module. In this section, we state each of them in details. For an illustrative walk-through example of ProofAug, refer to Figure 1.
@@ 3.1 Find the Maximal Compatible Semi-Proof
The FindMCSP routine is described in Algorithm 1. Firstly, $y_{f}^{0}$ is parsed into a sequence of proof steps $\mathbf{a}$. Then we create a pointer $i$ indicating the current position and an array $\mathbf{s}$ indicating the states before the $i$th step. In each iteration, we first try to proceed with current proof step $\mathbf{a}[i]$ from the current state $\mathbf{s}[i]$ in the ITP. If this step succeeds, we continue to the next proof step; otherwise, according to the proof mode of $\mathbf{s}[i]$, we take different strategies: if $\mathbf{s}[i].\mathrm{mode}$ is proof(prove), we replace $\mathbf{a}[i]$ with sorry; otherwise, we find the innermost proofqed block of $y_{f}^{0}$ that contains the $i$-th step and substitute the whole block with a single sorry step, then move the pointer accordingly. The algorithm terminates when a failed step occurs outmost or all steps pass the check of ITP.
@@ 3.2 Proof Augmentation
The goal of ProofAug is to determine whether there exists a compatible semi-proof that can be completed into a valid proof using ATPs or built-in heuristic proof methods to fill the sorry gaps. Instead of first finding out all compatible semi-proofs and trying them one-by-one, we should design a more efficient way that makes use of the overlap among these semi-proofs to avoid waste of computation resources spent on calling ATPs. The resulted proof augmentation algorithm is shown in Algorithm 2 (Ignore the ERP module for now).
@@
For a concrete example of ProofAug inducing a proof from a failed initial proof proposal, refer to Appendix H.
@@ 3.3 Efficient Recursive Proving
In order to achieve a more sample efficient and fine-tuning free recursive proving method, we design our algorithm in a modularize way under the unified view described in Section 2.1. Firstly, for the prompt design module that determines the expected output content, we borrow the proof completion module of RMaxTS to replace the proof-generation step of POETRY; then, for the response processing module, we perform proof structure analysis as in the proof augmentation procedure and try <ATP>monospace-$<ATP>|$> for the failed proofqed blocks or by clauses. A new node is added if the <ATP>monospace-$<ATP>|$> method fails to solve the corresponding conjecture. For the state selection strategy, we follow the best-first strategy, but the priority is modified from the cumulative log-probability to the end position of the failed proof block, i.e., if one node is an ancestor of another, we prioritize the descendant; otherwise, we prioritize the one that appears first in the semi-proof.
In practice, to simplify the algorithm and suppress the sampling cost wasted in low-level details, we add new nodes and try <ATP>monospace-$<ATP>|$> for failed proofqed blocks in the first iteration only. Such simplifications makes the above procedure become a plug-and-play ERP module over our ProofAug, as shown in Algorithm 2. This coincidence can be interpreted by that each conjecture naturally inherits one proof try from the original proof proposal. In other words, the ProofAug without ERP module is already implicitly a recursive proving method, in the sense that each compatible semi-proof exactly corresponds to one proof tree during the search process of POETRY if the the two methods share the same final proof.
For an illustrative example of how the ERP module works in ProofAug, refer to Figure 1.
@@
Cumulative deepseek-math-7b-base 1400 61.9%
Cumulative + Dataset Curation deepseek-math-7b-base 2100 66.0%

### v2

@@ Abstract
The synergy between deep learning models and traditional automation tools, such as built-in tactics of the proof assistant and off-the-shelf automated theorem provers, plays a crucial role in developing robust and efficient neural theorem provers (NTPs). However, for proof synthesis with LLMs, previous work applies automation tools either only when explicitly invoked by the model or at a single granularity level, failing to fully exploit their power. To solve this issue, we propose ProofAug, a procedure that equips LLMs with automation methods at various granularities through fine-grained structure analysis of model-generated proof proposals. ProofAug also serves as a versatile plug-and-play module that seamlessly integrates with any tree-search algorithm, enabling our construction of an efficient recursive proving (ERP) module to further enhance performance. The superiority of our method is validated on the miniF2F benchmark using the open-source deepseek-math-7b-base model and the Isabelle proof assistant. Notably, by additionally employing a mixed prompting strategy, we achieve a cumulative pass rate of 66.0% after curation of the dataset (61.9% for the original version) with at most 2100 queries to the model per problem (In contrast, the previous SOTA in Isabelle, Subgoal-XL (Zhao et al., 2024), only achieves 56.1% using 16384 queries per problem). We also implement a Lean 4 version of ProofAug that can improve the pass@1 performance of Kimina-Prover-Preview-Distill-1.5B from 44.3% to 50.4% on miniF2F-test. Our code is available at https://github.com/haoxiongliu/ProofAug.
Machine Learning, ICML
@@ 1 Introduction
Automated theorem proving is a field that not only appeals to mathematicians looking for efficient formalization and proof automation for mathematical theorems, but also finds significant applications in real-world industries such as integrated circuit design and software engineering (Bibel, 2013; Schumann, 2013). Early applications of machine learning methods on automated theorem proving focus on tasks such as premise selection (Irving et al., 2016) and proof search (Loos et al., 2017) in SMT solvers or resolution/superposition provers, often interleaved with handcrafted heuristics. Recently, rapid progress in mathematical reasoning with large language models (LLMs) has awoken a flourishing interest for directly synthesizing formal proofs using generative language models.
@@
In literature, there are primarily two paradigms for direct proof synthesis: the proof-step generation paradigm and the whole-proof generation paradigm, each corresponding to a distinct proof style. For proof assistants where proofs are mainly in the procedural style (or tactic-style), such as MetaMath (Megill & Wheeler, 2019) and HOL Light (Harrison, 2009), previous work adopts the proof-step generation paradigm (Polu & Sutskever, 2020; Lample et al., 2022; Han et al., 2022; Polu et al., 2022), where in each iteration the model is requested to generate one tactic to be applied in the interactive theorem prover (ITP) [^1] We use the terms ‘interactive theorem prover’ and ‘proof assistant’ interchangeably with a nuance: the former emphasizes the formal system aspects, while the latter highlights its role in assisting users.. Then the ITP provides information on the updated proof state, which sometimes could be indispensable for people to finish the remaining proof. Proof-step generation methods typically use tree-search algorithms to guide the proof process, inspired by the success of AlphaGo (Silver et al., 2016). In contrast, for proof systems where declarative style proofs are common, such as Isabelle and Lean 4, computation-efficient whole-proof generation methods become possible. However, the efficiency in this paradigm is not free. Any incorrect application of a proof method that bridges the gap between two proof terms (or goals/conjectures) will cause the proof to fail to pass the ITP verification check.
@@
To avoid this issue, the Draft, Sketch, and Prove (DSP) framework (Jiang et al., 2023) proposes to let the model first write a formal proof sketch (Wiedijk, 2004) according to an informal proof draft, then offload the proofs of intermediate conjectures to off-the-shelf ATPs [^2] For simplicity, we follow the practice of Jiang et al. (2023) to use the abbreviation ATP to refer to any automated method, including SMT solvers and what traditionally known as automated theorem provers. and heuristic proof methods. Many subsequent works (Wang et al., 2023a; Zhao et al., 2023; Zheng et al., 2024; Zhao et al., 2024) also follow the practice of DSP when designing their few-shot prompts: The proof methods in the original proofs of the demonstration examples are replaced with a placeholder, which signifies that ATPs are responsible for completing the task. We observe that while this approach can significantly enhance performance compared to direct full-proof generation, it may also result in proof failures in unintended ways:
@@
  • (Complicated Draft) Due to common inconsistencies between informal proof and formal system, there are conjectures that despite being easy to be solved by ATPs, humans need to write complicated informal proofs for them. The unnecessary intermediate steps could be translated into formal conjectures that are inappropriate in syntax or even harder than the original theorem.
@@
Figure 1: Illustration example of ProofAug. Each box in the flowchart corresponds to a proof step. (a) The initial proof encounters an error when proving Conjecture 1-0. (b) We replace the original proof of Conjecture 1-0 with a sorry and continue, until a syntactic error occurs at Conjecture 1-1. (c) The proof of Conjecture 1 is replaced by sorry and we obtain a semi-proof with two sorrys. (d) Automation tools are called to prove Conjecture 0 and Conjecture 1, but failing at the latter one. (e) The ERP module (introduced in Section 3.3) sends a completion request at Conjecture 1. (f) The generated ERP proof still fails at Conjecture 1. (g) We finally resort to a more coarse level of proof and successfully find a proof with automation tools. Refer to Algorithm 2 for a pseudo-code description of ProofAug.
@@
In this paper, we propose to address this dilemma through performing fine-grained structural analysis on the proof proposals generated by LLMs, based on the observation that regardless of correctness, full proofs contain rich structural information from which valid semi-proofs [^4] By semi-proofs, we refer to proofs that possibly contain symbols indicating pending proof, such as sorry in Isabelle and Lean. of varying granularities can be derived. Specifically, given a theorem statement, instead of instructing the model to generate a proof sketch as previous work does, we choose to start from generating a full proof to suppress the ‘Hard Conjectures’ issue in the first place. For the generated proof proposal, we first find its corresponding maximal compatible semi-proof (MCSP) that maintains most structural information yet can pass the check of ITP. Then, departing from it, we recursively resort to a more coarse semi-proof whenever ATPs fail to fill in some gap in the current one. This procedure solves the ‘Complicated Draft’ issue, thus significantly improving the sample efficiency. We call this procedure ProofAug. Figure 1 shows an example flow of our method.
@@
We also demonstrate that the application of proof structure analysis is not restricted to single-pass generation. In fact, ProofAug can be incorporated into any tree-search algorithm to expand the candidate node set at each iteration. This will become evident in our unified view of theorem proving introduced in Section 2.1. In Section 3.3, we illustrate its application in constructing a sample-efficient recursive proving (Wang et al., 2024) module on top of ProofAug.
@@
We list our contributions with corresponding key evaluation results below. All experiments use the open-source deepseek-math-7b-base model (Shao et al., 2024) if not otherwise specified.
@@
  • We propose a versatile plug-and-play procedure ProofAug that can significantly enhance the performance of neural theorem provers, based on fine-grained proof structure analysis. On miniF2F-test [^5] We use ‘miniF2F-test’ to refer to the miniF2F test split Isabelle data included in the code repository of Jiang et al. (2023). See Section 4.1 for a detailed discussion on the versions of miniF2F. (Zheng et al., 2021), ProofAug achieves a pass rate of 52.5% with no more than 100 queries to the model. Ablation results show that when compared with the DSP baseline, the improvement brought by ProofAug implementation for Isabelle is 7.8%, 4.1% and 3.3% under sample budgets of 1, 10 and 100, respectively.
@@
  • ProofAug can be incorporated into any tree-search algorithms under our unified view of existing generative LM based theorem proving methods. We showcase that when integrated with our Efficient Recursive Proving (ERP) module, 0-shot ProofAug performance improves from 54.5% to 56.1% with a limit of 500 queries per problem. For comparison, RMaxTS (Xin et al., 2024b) only shows an advantage of $<0.2\%$ over naive single-pass retrying under a sample budget of 3200.
@@
  • Moreover, using a mixture of strategies with different prompting methods, we achieve a total pass rate of 61.9% within 1700 model queries. After curating the miniF2F-test data, we manage to prove 66.0% of the theorems with no more than 2100 requests, setting a new state-of-the-art across methods using the Isabelle proof assistant.
  • We also build a Lean implementation of ProofAug and shows that the pass@1 accuracy of Kimina-Prover-Preview-Distill-1.5B on miniF2F-test can be improved from 44.3% to 50.4%, exhibiting the generality of ProofAug for various proof systems.
@@ 2.1 A Unified View of Theorem Proving with Generative Language Models
[…]
@@
For any string $x\in\Sigma^{*}$, we use $s_{x}$ to denote $T(s_{0},x)$, where we have lifted the notation $T$ to express consecutive state transitions, i.e., for $s\in\mathcal{S}$,
$T(s,x):=T(\cdot,\mathbf{a}[n])\circ T(\cdot,\mathbf{a}[n-1])\circ\cdots\circ T(\cdot,\mathbf{a}[1])(s)$
if $\mathrm{Parse}(x)=(\mathbf{a}[1],\mathbf{a}[2],\cdots,\mathbf{a}[n])$.
@@
2.2 The Isabelle/Isar Proof Language
Although the basic idea of our method in general fits any ITP, the implementation details vary depending on specific features of the proof assistant. In this paper, we use Isabelle/Isar (Wenzel et al., 2004) as an example to introduce the spirit of ProofAug and how to implement it in a specific proof system. [^7] See Appendix C for a detailed explanation of why our method is generic across different systems and a brief summary of our Lean implementation. In this section, we introduce some basic knowledge and unique features that ease the implementation of our method in Isar.
@@ 3 Method
Algorithm 2 describes our method, ProofAug. It mainly consists of three parts: finding the Maximal Compatible Semi-Proof (MCSP), proof augmentation, and an optional Efficient Recursive Proving (ERP) module. In this section, we state each of them in detail. For an illustrative walk-through example of ProofAug, refer to Figure 1.
@@ 3.1 Find the Maximal Compatible Semi-Proof
The FindMCSP routine is described in Algorithm 1. Firstly, $y_{f}^{0}$ is parsed into a sequence of proof steps $\mathbf{a}$. Then we create a pointer $i$ indicating the current position and an array $\mathbf{s}$ indicating the states before the $i$th step. In each iteration, we first try to proceed the current proof step $\mathbf{a}[i]$ from the current state $\mathbf{s}[i]$ in the ITP. If this step succeeds, we continue to the next proof step; otherwise, according to the proof mode of $\mathbf{s}[i]$, we take different strategies: if $\mathbf{s}[i].\mathrm{mode}$ is proof(prove), we replace $\mathbf{a}[i]$ with sorry; otherwise, we find the innermost proofqed block of $y_{f}^{0}$ that contains the $i$-th step and substitute the whole block with a single sorry step, then move the pointer accordingly. The algorithm terminates when a failed step occurs outmost or all steps pass the check of the ITP.
@@ 3.2 Proof Augmentation
The goal of ProofAug is to determine whether there exists a compatible semi-proof that can be completed into a valid proof using ATPs or built-in heuristic proof methods to fill the sorry gaps. Instead of first finding out all compatible semi-proofs and trying them sequentially, we should design a more efficient way that makes use of the overlap among these semi-proofs to avoid waste of computation resources spent on calling ATPs. The resulting proof augmentation algorithm is shown in Algorithm 2 (Ignore the ERP module for now).
@@
For a concrete example of ProofAug inducing a proof from a failed initial proof proposal, refer to Appendix I.
@@ 3.3 Efficient Recursive Proving
In order to achieve a more sample efficient and fine-tuning free recursive proving method, we design our algorithm in a modularized way under the unified view described in Section 2.1. Firstly, for the prompt design module that determines the expected output content, we borrow the proof completion module of RMaxTS to replace the proof-generation step of POETRY; then, for the response processing module, we perform proof structure analysis as in the proof augmentation procedure and try <ATP>monospace-$<ATP>|$> for the failed proofqed blocks or by clauses. A new node is added if the <ATP>monospace-$<ATP>|$> method fails to solve the corresponding conjecture. For the state selection strategy, we follow the best-first strategy, but the priority is modified from the cumulative log-probability to the end position of the failed proof block, i.e., if one node is an ancestor of another, we prioritize the descendant; otherwise, we prioritize the one that appears first in the semi-proof.
In practice, to simplify the algorithm and suppress the sampling cost wasted in low-level details, we only try <ATP>monospace-$<ATP>|$> and add new nodes for failed conjectures belonging to the original MCSP. Such simplifications make the above procedure become a plug-and-play ERP module over our ProofAug, as shown in Algorithm 2. This coincidence can be interpreted as that each conjecture naturally inherits one proof try from the original proof proposal. In other words, the ProofAug without ERP module is already implicitly a recursive proving method, in the sense that each compatible semi-proof exactly corresponds to one proof tree during the search process of POETRY if the two methods share the same final proof.
For an illustrative example of how the ERP module works in ProofAug, one can refer to Figure 1.
@@
ProofAug + Mixed Strategy deepseek-math-7b-base 1400 61.9%
ProofAug + Mix. + Dataset Curation deepseek-math-7b-base 2100 66.0%